OpenStudy (anonymous):

Fan&Medal!!!! Solve for x x2 - 3|x - 2| - 4x = - 6

OpenStudy (anonymous):

start by isolating |x-2| to the left hand side

OpenStudy (anonymous):

@Chris01 are you still here?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

can you bring |x-2| to one side, and all the other terms to the other side?

OpenStudy (anonymous):

sorry i'm confused I don't know how to do that part

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

x^2 - 3|x - 2| - 4x = - 6 let's first subtract x^2 from both sides so we'll have x^2 - x^2 -3| x-2| -4x = -6 -x^2 = -3| x-2| -4x = -6 -x^2

OpenStudy (anonymous):

then we add 4x to both sides to cancel out the -4x as well so we get -3| x-2| = -6 -x^2 +4x

OpenStudy (anonymous):

now we divide both sides by -3

OpenStudy (anonymous):

| x-2| = (-6 -x^2 +4x) / -3 \[\left| x-2 \right| = 2*\frac{ x^2 }{ 3 }-\frac{ 4x }{ 3 }\]

OpenStudy (anonymous):

now, once we finally have |x-2| isolated we can eliminate the absolute stripes. we do this by making two versions of the equation, one positive and one negative

OpenStudy (anonymous):

\( \Large \left| x-2 \right| = 2+\frac{ x^2 }{ 3 }-\frac{ 4x }{ 3 } \) becomes \( \Large x-2 = 2+\frac{ x^2 }{ 3 }-\frac{ 4x }{ 3 } \) and \( \Large x-2 =- 2-\frac{ x^2 }{ 3 }+\frac{ 4x }{ 3 } \)

OpenStudy (anonymous):

now we solve both of these for x

OpenStudy (anonymous):

starting with \( \Large x-2 = 2+\frac{ x^2 }{ 3 }-\frac{ 4x }{ 3 }\) probably best to multiply both sides by 3 to get rid of fractions \( \Large 3x-6 = 6+x^2-4x\) now get everything to one side \( \Large 0 = 6+x^2-4x-3x+6\) simplify \( \Large x^2-7x+12=0\) factorize \( \Large (x-4)(x-3)=0\) x=3 or x=4 that's already two of the 4 solutions if you solve the other part you get the other two solutions

OpenStudy (anonymous):

thanks man!!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

to clarify, the other two solutions are in \(\Large x-2 =- 2-\frac{ x^2 }{ 3 }+\frac{ 4x }{ 3 }\) (the 'negative' part of the equation)